how much of an alloy that is 30 copper should be mixed with 100 ounces that is 70% copper in order to get an alloy that is 50% copper
Answer by TimothyLamb(4379)

a = weight of 30% copper allow
w = weight of final mixture
---
30a + 70(100) = 50w
w = 100 + a
---
30a + 7000 = 50w
30a + 7000 = 50(100 + a)
30a + 7000 = 5000 + 50a
20a = 2000
a = 100
---
answer:
mix in 100 ounces of 30% copper alloy
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php